OpenStudy (madmerc):

What is the simplified form of 4 sqrt x^2 times 4 sqrt x^2? ( i think it looks like 4√ײ times 4√ײ ) the answer choices are x^1/2 x x^2 √x

OpenStudy (puppylover784):

Need help

OpenStudy (puppylover784):

First Tengo que ir a caca .

Nnesha (nnesha):

is it \[\huge\rm \sqrt[4]{x^2} * \sqrt[4]{x^2}\] 4 as an index ?

OpenStudy (madmerc):

yes thats it! i couldn't find the thing to do that

Nnesha (nnesha):

alright apply the exponent rule you can convert radical to an exponent form \[\huge\rm \sqrt[\color{Red}{n}]{x^\color{blue}{m}}=x^\frac{\color{blue}{ m }}{ \color{ReD}{n} }\]index becomes denominator of fraction

Nnesha (nnesha):

the*

OpenStudy (madmerc):

so it's x^2/4 ?

Nnesha (nnesha):

yes convert both radical to an exponent fform \[\huge\rm x^\frac{ 2 }{ 4 }*x^\frac{2}{4}\]

Nnesha (nnesha):

and when we multiply same bases we should `add` exponents \[\huge\rm x^m \times x^n=x^{m+n}\]

OpenStudy (madmerc):

so it's 3x?

OpenStudy (madmerc):

I don't understand

Nnesha (nnesha):

how did you get 3 ?

OpenStudy (madmerc):

because 2+2 is 4 and 4 over 4 is 1 there are 2 x's so 3x? idk i'm bad at math

Nnesha (nnesha):

alright first we converted radical to an exponent form right \[\huge\rm \color{ReD}{ \sqrt[4]{x^2}} * \color{blue}{\sqrt[4]{x^2}}\] \[\huge\rm\color{Red}{x^\frac{ 2 }{ 4 }}*\color{blue}{x^\frac{2}{4}}\] you can reduce the fraction first 2/4 and then add the exponents

Nnesha (nnesha):

here is an example \[\huge\rm x^4 *x^4 =x^{4+4}=x^{8}\] we end up with only one base

Nnesha (nnesha):

x^8 is shorthand forx^4 * x^4

Nnesha (nnesha):

so when we combine them we should write it in terms of single base \[\huge\rm\color{Red}{x^\frac{ 2 }{ 4 }}*\color{blue}{x^\frac{2}{4}}\] \[\huge\rm x^{\frac{2}{4}+\frac{2}{4}}\]

OpenStudy (madmerc):

so it's x^2 ?

Nnesha (nnesha):

hmm how did you get 2 ?

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @madmerc because 2+2 is 4 and 4 over 4 is 1 there are 2 x's so 3x? idk i'm bad at math \(\color{blue}{\text{End of Quote}}\) right 2+2 = 4 and 4/4 =1 that's correct but there is only one x

OpenStudy (madmerc):

so is it just x as the answer?

Nnesha (nnesha):

right which is same as x^1

OpenStudy (madmerc):

thank you so much

Nnesha (nnesha):

np :=))